# Gas contamination and mitigation in a 100 m3 / 10 bar argon TPC with optical readout: a viability study

### v1

@@ Abstract
Gaseous Optical Time Projection Chambers (OTPCs) aimed at Neutrino Physics and Rare Event Searches will likely exceed the tonne scale during the next decade. This will make their performance sensitive to gas contamination levels as low as 100 ppb, that is challenging at room temperature due to outgassing from structural materials. In this work we discuss gas distribution and impurity mitigation on a 5 m-length/5 m-diameter 10 bar TPC filled with Ar/CF4 admixed at 99/1 per volume (1.75 tonne), loaded with technical plastics in order to enhance light collection and readout. Different distributor topologies, outgassing and flow rates are discussed. Specifically, our work is aimed at illustrating the conceptual viability of the optical readout of ND-GAr’s TPC (within the DUNE Near-Detector complex), in terms of material compatibility. For our proposal, with perforated distributors aligned with the electric field, and under realistic assumptions, the concentration of contaminants can be controlled within a week from chamber filling. In the case of N2, injection of fresh gas at %-level seems to represent the safest strategy to keep the concentration within operability limits.
@@
Gaseous detectors; Time projection Chambers; Outgassing; CFD simulations; Optical Readout
1 Introduction
Gas distribution and purification, together with material selection, are crucial aspects for the operation of gaseous instrumentation (e.g., [1]). With future large-volume detectors poised to reach the tonne scale in gas mass for the first time, the associated challenges will increase, particularly in the form of time projection chambers, or TPCs [2, 3, 4]. As an example, the DUNE experiment (dedicated to the study of neutrino oscillations) is considering the construction of a nearly 2 tonne argon-rich TPC that would operate at 10 bar, with 5 m diameter and 5 m drift (as part of the ND-GAr detector [5, 6]); the NEXT experiment (targeting the discovery of $\beta\beta 0\nu$) aims at a bit over 1 tonne of 136Xe in a 15 bar TPC of 2.6 m (diameter)×${}\times 2.6$ m (drift) [7]; CYGNUS-1000 (whose goal is the reconstruction of the direction of galactic WIMP dark matter) plans operation under He/SF6(CF4) at around atmospheric pressure in a 1000 m3 detector, totaling 3 tonne [8]. The need for such unprecedented masses of gas is driven by the unique capability of this phase to resolve the details of the interactions involving low-energy particle tracks, provided a sufficient interaction/decay probability can be guaranteed in the first place.
There are, as well, contemporary elements that further complicate the design at this new scale: some of the gases considered (e.g., SF6, CF4) represent a threat to the environment, [^1] Available estimates of the global warming potential (GWP) exceed 20000 for SF6 and 5000 for CF4, [9]. while others (e.g., 136Xe) are too precious to be vented. This enforces operation in closed-loop, thereby binding the system performance to the power of the compressors and purification systems, in competition with the outgassing rate from internal elements. When high gas mass is achieved through pressurization, the attachment probability (electron lifetime) is increased (decreased) as $\sim{P}^{2}$ [10] and both charge collection and mm-level space resolutions can be compromised even in the presence of tiny concentrations of contaminants. As shown in this work, the sensitivity of the drift velocity to impurities, combined with experimental values of the attachment coefficient, set the tolerable H2O and O2 levels at the 0.1-1 parts per million (ppm) scale for $\mathcal{O}(10~{}$m) drift. This is at least one order of magnitude below the purity requirements of the ALICE TPC, the largest gaseous TPC to date [11].
Simulations based on computational fluid dynamics (CFD) are nowadays possible in a variety of systems. As they provide critical information on the flow characteristics and the admixing of species, they can reduce the investment in physical prototypes and help anticipating problems. CFD simulations of gaseous detectors usually focus on characteristics of the gas flow such as gas stagnancy (‘pockets’) and the impact of temperature gradients, since they modify the detector response locally [12]. However, not much has been done to address the role of contaminants, to the best of the authors’ knowledge. Modeling the dynamics of contaminants in fluid detectors is not alien to modern technologies, and it has been successfully done for instance for liquid argon TPCs based on a set of differential equations describing each of the elements in the system [13].
In this work we study gas distribution and impurity mitigation in the context of future large volume/large mass TPCs intended for experiments in the field of Neutrino Physics and Rare Event Searches. We discuss, in particular, a scenario where a massive use of technical plastics would be made in order to enhance the optical response of a TPC aiming at the study of neutrino interactions, following the conceptual proposal outlined in [14]. A distributor geometry based on axial rods has been chosen as the baseline, although other gas distribution concepts are possible (e.g. through forced axial flow). Axial distribution rods have the advantage of being a priori amenable to any field-cage design and little dependent on the end-cap details.
The structure of this document is as follows. First, section 2 discusses the main inputs used in the CFD simulations: (2.1) the influence of contaminants on the TPC response, obtained from existing data or electron-transport simulations; (2.2) outgassing rates of technical plastics; and (2.3) gas distribution scheme. Section 3 deals with the optimization of the geometry of the distributor rods (3.1), and the introduction of the CFD simulation framework (3.2). Section 4 presents our main results, chiefly the spatial and temporal distributions of common outgassing species in the TPC for different distribution schemes, species and conditions. Finally, we end with our conclusions in section 5.
2 Inputs
We adopt a reference geometry of a 5 m-length/5 m-diameter cylindrical TPC, as that currently under consideration for DUNE’s ND-GAr ([5, 6]), intermediate in size with respect to [7] and [8]. The recently introduced possibility of performing ‘Full3D’ optical tracking in Ar/CF4 (99/1) [14, 15, 16, 17], involves the use of large amounts of reflective (Teflon) and transparent (PMMA) plastics, to enhance the optical response of the detector. That extreme situation (in terms of system outgassing) is discussed in the following. A scheme of the system, together with a conceptual gas distribution scheme, can be found in Fig. 1.
2.1 Impact of impurities on TPC response
Contaminants such as H2O, O2 and N2 will be considered, since they are predominantly found in the outgassing of materials with low vapour pressure (i.e., barring most epoxies and adhesives from detector construction). We neglect contamination from exogenous impurities arising for instance from residual gas in the bottles, that might be pernicious at concentrations much below the ones discussed in this work [18]. Contaminants such as CO2 and CO can be also found, e.g. for technical plastics [19], at outgassing rates similar to those of the aforementioned species. An upper limit to their impact on the chamber performance can be obtained from the analysis made in this section for the case of H2O, since the latter displays one of the strongest electron-cooling powers of the molecules discussed: this means, namely, a larger distortion of the drift of the primary ionization electrons compared to the purified gas. The main mechanisms through which impurities can limit the performance of a TPC are: i) the capture of the drifting electrons (attachment) and ii) the alteration of the drift velocity and the ensuing track distortions, affecting momentum reconstruction. For TPCs relying on scintillation, as discussed here, iii) modifications to the scintillation probability and absorption can occur too.
Reference values for the contaminant thresholds as used in this paper can be determined by examining the effects in attachment, drift velocity and light quenching and absorption, as follows:
  • Attachment: of upmost importance at high pressure is the attachment probability per unit length for electrons drifting in the presence of oxygen ($\eta$), or its related magnitude: the electron lifetime ($\tau_{e}$). For argon-based mixtures those may be parameterized, after [10], as:
    $\eta$ $=$ $\frac{1}{v_{d}}\cdot\left(1+\frac{f_{H_{2}O}}{1000}\right)\cdot P^{2}\cdot f_{O_{2}}\cdot\mathcal{A}_{\text{gas}},$ (1)
    $\tau_{e}$ $=$ $\frac{1}{\eta\cdot v_{d}},$ (2)
    where the magnitude $\mathcal{A}_{\text{gas}}$ is gas-mixture dependent, $f_{O_{2}(H_{2}O)}$ represents the volume concentration of O2 (H2O), $v_{d}$ is the electron drift velocity and $P$ is the gas pressure. The $P^{2}$-dependence reflects the fact that the dominant attachment process takes place in two steps (Bloch-Bradbury-Herzengber mechanism, BBH, [20, 21]):
    $e^{-}+\textnormal{O}_{2}$$\rightarrow$ $\textnormal{O}_{2}^{-,*},$ (3)
    $_{2}^{-,*}+\textnormal{O}_{2}~{}(\textnormal{H}_{2}\textnormal{O})$ O ) $\rightarrow$ $\textnormal{O}_{2}^{-}+\textnormal{O}_{2}^{*}~{}(\textnormal{H}_{2}\textnormal{O}^{*}).$ (4)
    Both $\mathcal{A}_{\text{gas}}$ and $v_{d}$ depend solely on the pressure-reduced electric field ($E/P$), therefore not showing separate $E$ and $P$ dependencies. [^2] This is a direct consequence of the fact that $E/P$ determines uniquely the electron energy distribution. Typical pressure-reduced drift fields in TPCs are found in the range of 10 – few 100’s V/cm/bar, expectedly below the dissociative-attachment thresholds for either CF4, O2 or H2O. This observation has been backed with the state-of-the-art Pyboltz transport code [22], thus anticipating BBH to be the dominant attachment mechanism in the conditions discussed in this work.
    The attachment coefficient, $\eta$, in Ar/CF4 (99/1) at $E/P=100$ V/cm/bar might be estimated starting from the experimental $\mathcal{A}_{\text{gas}}$ value in case of the classical Ar/CH4 mixture at 90/10 (‘P10’). The compilation in [10], performed up to 4 bar, indicates that $\mathcal{A}_{\text{gas}}=15.1\pm 1.5{\mu}$s-1, and $v_{d}=5.36$ cm/$\mu$s. Extrapolating to 10 bar following eq. 1, this value would correspond to 10% charge loss over 5 m drift for O2 concentrations of 0.7 ppm. The same work reports a linear decrease of $\mathcal{A}_{\text{gas}}$ with the decrease of the concentration of the molecular additive and an increase with the size of the molecule (e.g., from CH4 to i-C4H10). The effect can be attributed to the dependence of the stabilization reaction 4 with the concentration of the molecular additive and with the molecule size. For Ar/i-C4H10 at 99/1, $\mathcal{A}_{\text{gas}}$ = $48.4\pm 3.7{\mu}$s-1 and $v_{d}=4.97$ cm/$\mu$s were measured at $E/P=100$ V/cm/bar, yielding a 10% charge drop over 5 m for 0.2 ppms of O2. Given that i-C4H10 and CF4 molecules are similar in size, a similar strength of reaction 4, and thus similar attachment rates, might be expected.
@@ 2.1 Impact of impurities on TPC response
    An $E/P$ of 100 V/cm/bar would correspond to an imposing 500 kV cathode bias on a 5 m drift distance at 10 bar. Based on the studies of breakdown voltage for Ar/CF4 (99/1) carried out in [23], as well as the availability of commercial feedthroughs, we will restrict our discussion to a cathode bias not greater than 250 kV ($E/P=50$ V/cm/bar). That would roughly double the value of $\mathcal{A}_{\text{gas}}$ compared to $E/P=100$ V/cm/bar, extrapolating from [10]. Based on this, we aim in this work for an O2 concentration value not larger than 0.1 ppm, so as to keep the charge losses well within 10%.
  • Modifications of the drift velocity: impurities, even in trace-amount, can substantially modify the drift velocity to the extent of limiting particle tracking, too. Fig. 2 provides the sensitivity of Ar/CH4 (90/10) to H2O (left) and N2 (right), as a function of the pressure-reduced electric field in the range 5-50 V/cm/bar. It represents the magnitude $\left|\frac{\Delta{v_{d}}}{v_{d}}\right|=\left|\frac{v_{d}(f=0)~{}-~{}v_{d}(f)}{v_{d}(f=0)}\right|$, obtained with Pyboltz at $P=10~{}$bar and $T=25~{}^{\circ}$. The figure shows how, as expected, the impact of H2O is stronger than that of N2 by at least one order of magnitude.
@@
    $\Delta{z}=L\left|\frac{\Delta v_{d}}{v_{d}}\right|\simeq\frac{L}{v_{d}(f=0)}\cdot{\left|\frac{dv_{d}}{df}\right|}\cdot f$ (5)
    with $L=5$ m being the drift distance. The linear approximation on the r.h.s. of eq. 5 allows extrapolation to low concentrations ($f$), for which the numerical precision of simulations is lost. Fig. 3 presents the maximal track distortion as a function of the concentration of impurities (from eq. 5) for two typical electric fields (200 V/cm, 400 V/cm) and three gas mixtures (Ar/CH4 at 90/10 and 98/2, and Ar/CF4 at 99/1). For reference, the 1 mm space-resolution landmark is indicated with a horizontal line. As seen, depending on the system requirements, the tolerable impurity levels range from 10 ppms of N2 (1 mm distortion) down to values as low as 0.05 ppms of H2O (0.1 mm distortion). The case of O2 is not shown since the contamination levels that are acceptable for tracking are well above those for H2O and, in most practical conditions, its presence will impact the electron lifetime much before it starts to be a concern for tracking. The tolerable concentrations of contaminants that will be assumed throughout this work will be 10 ppm for N2 and 1 ppm for H2O (limited by distortions in tracking at mm-level) and 0.1 ppm for O2 (limited by attachment). They are compiled in table 1.
@@
    $P_{Q}=1-P_{\text{scin}}\simeq 1-\frac{1}{k\cdot P\cdot f\cdot\tau_{3}}$ (6)
    Here $P_{\text{scin}}$ is the scintillation probability, $k$ the quenching rate in units of bar-1, and $f$ is the fraction of contaminant. Measurements of the scintillation quenching in [29] point to 0.2 ppm as the concentration at which CO2 would cause 10%-level light losses, and about 20 ppm for N2 (at 10 bar pressure).
    The experimental values obtained above, for the maximum concentrations of impurities that are compatible with argon scintillation, are at the same level (or even below) than those obtained from the analysis of the charge properties in regard to attachment and drift velocity. The situation for Ar/CF4 (99/1) is expected to be much more favourable, though. First, about 90% of the usable scintillation lies in the near-UV and visible range (200-800 nm) in that case [31, 16]. Considering the absorption coefficient for H2O at the low-end of such wavelength range, a value of $\Pi=10^{-2}$ cm-1 bar-1 is expected [28]. As a result, 10%-level light losses would require of H2O concentrations as high as 200 ppms (being water the most opaque contaminant, a priori). Regarding light quenching, it has been measured in [32] for pure CF4 at 1 bar to be 50% under a N2 contamination of 4%. Assuming a quenching relation of the type 6, 4000 ppms of N2 would be needed to cause a 10% light loss. The factor 200 above the values obtained for pure argon matches well the expectation in eq. 6 resulting from the much shorter lifetime of CF4 states, in the range 10-20 ns [15]. Following the analogy with the argon case further, a plausible estimate of the maximum amount of (more reactive) molecules such as CO2 and H2O would be 40 ppms, greatly exceeding the upper bounds obtained from charged transport.

### v2

@@ Abstract
Gaseous Optical Time Projection Chambers (TPCs) aimed at Neutrino Physics and Rare Event Searches will likely exceed the tonne scale during the next decade. This will make their performance sensitive to gas contamination levels as low as 100 ppb, which is challenging at room temperature due to outgassing from structural materials. In this work we discuss gas distribution and impurity mitigation in a 5 m-length/5 m-diameter 10 bar TPC filled with Ar/CF4 admixed at 99/1 per volume (1.75 tonne), loaded with technical plastics in order to enhance light collection and readout. Different distributor topologies, outgassing and flow rates are discussed. Specifically, our work is aimed at illustrating the conceptual viability of the optical readout of ND-GAr’s (Near Detector - Gaseous Argon) TPC (within the DUNE Near-Detector complex), in terms of material compatibility. For our proposal, with perforated distributors aligned with the electric field, and under realistic assumptions, the concentration of contaminants can be controlled within a week after chamber filling. In the case of N2, injection of fresh gas at %-level seems to represent the safest strategy to keep the concentration within operability limits.
1 Introduction
Gas distribution and purification, together with material selection, are crucial aspects for the operation of gaseous detectors (e.g., [1]). With future large-volume detectors (in particular time projection chambers, or TPCs) poised to reach the tonne scale in gas mass for the first time, the associated challenges will increase [2, 3, 4]. DUNE, NEXT and CYGNUS-1000 are several examples where the use of very large gaseous TPCs is foreseen. The DUNE experiment (dedicated to the study of neutrino oscillations) is considering the construction of a nearly 2 tonne argon-rich TPC. This TPC would operate with gaseous argon (GAr) at 10 bar, with 5 m diameter and 5 m drift as part of the Near Detector (ND) complex – such a detector system is known as ND-GAr [5, 6]. The NEXT experiment (targeting the discovery of $\beta\beta 0\nu$) aims at a bit over 1 tonne of 136Xe in a 15 bar TPC of 2.6 m (diameter)×${}\times 2.6$ m (drift) [7]. CYGNUS-1000 (whose goal is the reconstruction of the direction of galactic WIMP dark matter) plans operation under He/SF6(CF4) at around atmospheric pressure in a 1000 m3 detector, totaling 3 tonne [8]. The need for such unprecedented masses of gas is driven by the unique capability of this phase to resolve the details of the interactions involving low-energy particle tracks, provided a sufficient interaction/decay probability can be guaranteed in the first place. Time Projection Chambers are based on the principle of drifting ionization electrons over distances of several meters onto a suitably pixelated anode: the detector obtains, additionally, the longitudinal position of these electrons from their arrival times, upon multiplication by the drift velocity $v_{d}$ (a property of the gas at a given electric field). Minute concentrations of gas pollutants can trap the electrons during their drift and, when the device is used for precise mm-level tracking, they can also cause modifications of the drift velocity at a mere $10^{-4}$ level, significantly distorting the reconstructed particle trajectories.
Gas distribution becomes complicated at the aforementioned scale, from practical considerations alone: some of the gases considered (e.g., SF6, CF4) represent a threat to the environment, [^1] Available estimates of the global warming potential (GWP) exceed 20000 for SF6 and 5000 for CF4 [14]. while others (e.g., 136Xe) are too precious to be vented. In fact, due to the sheer amount of gas mass, it can be argued that virtually any open loop scheme will become impractical and uneconomical for tonne-scale TPCs, compared to a closed-loop implementation. Closed-loop operation binds the system performance to the power of the compressors and purification systems, in competition with the outgassing rate from internal elements. When high gas mass is achieved through pressurization, the attachment probability (electron lifetime) is increased (decreased) as $\sim{P}^{2}$ [9] and both charge collection and mm-level space resolutions can be compromised even in the presence of tiny concentrations of contaminants. As shown in this work, the sensitivity of the drift velocity to impurities, combined with experimental values of the attachment coefficient, sets the tolerable H2O and O2 levels at the 0.1-1 parts per million (ppm) scale for $\mathcal{O}(10~{}$m) drift. This is at least one order of magnitude below the purity requirements of the ALICE TPC, the largest gaseous TPC to date [10].
Simulations based on computational fluid dynamics (CFD) are nowadays possible in a variety of systems. As they provide critical information on the flow characteristics and the admixing of species, they can reduce the investment in physical prototypes and help anticipate problems. CFD simulations of gaseous detectors usually focus on characteristics of the gas flow such as gas stagnancy (‘pockets’) and the impact of temperature gradients, since they modify the detector response locally [11]. However, not much has been done to address the role of contaminants, to the best of the authors’ knowledge. Modeling the dynamics of contaminants in fluid detectors is not alien to modern technologies, and it has been successfully done for instance for liquid argon TPCs based on a set of differential equations describing each of the elements in the system [12].
In this work we study gas distribution and impurity mitigation in the context of future large volume/large mass TPCs intended for experiments in the field of Neutrino Physics and Rare Event Searches. We discuss, in particular, a scenario where a massive use of technical plastics would be made in order to enhance the optical response of a TPC aiming at the study of neutrino interactions, following the conceptual proposal outlined in [13].
The structure of this document is as follows. First, section 2 discusses the main inputs used in the CFD simulations: (2.1) the influence of contaminants on the TPC response, obtained from existing data or electron-transport simulations; (2.2) outgassing rates of technical plastics; and (2.3) the gas distribution scheme. Section 3 deals with the optimization of the geometry of the distributor rods (3.1), and the introduction of the CFD simulation framework (3.2). Section 4 presents our main results, chiefly the spatial and temporal distributions of common outgassing species in the TPC for different distribution schemes, species and conditions. Finally, we end with our conclusions in section 5.
2 Inputs
We adopt as a reference geometry a 5 m-length/5 m-diameter cylindrical TPC, similar to the one under consideration for DUNE’s ND-GAr [5, 6], intermediate in size between [7] and [8]. The recently introduced possibility of performing ‘Full3D’ optical tracking in Ar/CF4 (99/1) [13, 15, 16, 17], involves the use of large amounts of reflective (Teflon/PTFE) and transparent (PMMA) plastics, to enhance the optical response of the detector. Teflon, in particular, can intrinsically achieve a reflectivity well above 95% in the visible region of the spectra, allowing a flatter light collection in the range 35% (anode)-65% (cathode) and up to ×$\times 7$ increased light collection for events coming from the distant-most (anode) region [13]. A diagram of the system and a conceptual gas distribution scheme are shown in Fig. 1.
It is assumed that the gas is continuously recirculated through compressors which, along with purifiers (getters), maintain the injected gas clean. Commercial ‘cold’ getters are capable of maintaining O2 and H2O concentrations at ppb-level per pass, while ignoring common molecular additives such as CH4 and CF4 (a getter fulfilling these requirements is model 905 in [18], for instance). In simulation, since purification levels are at least two orders of magnitude below harmful concentrations, we assume the injected gas is always 100% pure. Conversely, getters that trap N2 require heating, which may be incompatible with high gas flows (due to accelerated cooling of the active materials of the getter) and molecular additives (due to higher reactivity compared to noble gases). For the nitrogen contamination studies performed here, we neglect the possibility of purification. [^2] An interesting discussion on the operation and performance of both types of getters can be found in [19].
2.1 Impact of impurities on TPC response
Contaminants such as H2O, O2 and N2 will be considered, as they predominantly originate from outgassing of low-vapor-pressure materials, excluding most epoxies and adhesives from detector construction. We disregard contamination from exogenous impurities, such as residual gas in bottles, which could be harmful at concentrations far below those considered here [20]. CO2 and CO contaminants can also originate from technical plastics, with outgassing rates comparable to the previously mentioned species [21]. Their impact on chamber performance can be upper-bounded using the H2O case, as it exhibits the strongest electron-cooling power among the discussed molecules. This results in a larger distortion of the drift of the primary ionization electrons compared to the purified gas. Impurities limit TPC performance mainly through: i) the capture of the drifting electrons (attachment) and ii) the alteration of the drift velocity and the ensuing track distortions, affecting momentum reconstruction. For TPCs that rely on scintillation, the one discussed here, iii) modifications to the scintillation probability and absorption can occur too. Reference values for the contaminant thresholds as used in this paper can be determined by examining the effects in attachment, drift velocity, and light quenching and absorption.
Attachment
Of utmost importance at high pressure is the attachment probability per unit length for electrons drifting in the presence of oxygen ($\eta$), or its related magnitude: the electron lifetime ($\tau_{e}$). For argon-based mixtures those may be parameterized, after [9], as:
$\eta$ $=$ $\frac{1}{v_{d}}\cdot\left(1+\frac{f_{H_{2}O}}{1000}\right)\cdot P^{2}\cdot f_{O_{2}}\cdot\mathcal{A}_{\text{gas}},$ (2.1)
$\tau_{e}$ $=$ $\frac{1}{\eta\cdot v_{d}},$ (2.2)
where the magnitude $\mathcal{A}_{\text{gas}}$ is gas-mixture dependent, $f_{O_{2}(H_{2}O)}$ represents the fraction of O2 (H2O) molecules, $v_{d}$ is the electron drift velocity and $P$ is the gas pressure. The $P^{2}$-dependence reflects the fact that the dominant attachment process takes place in two steps (Bloch-Bradbury-Herzengber mechanism, BBH, [22, 23]):
$e^{-}+\textnormal{O}_{2}$$\rightarrow$ $\textnormal{O}_{2}^{-,*},$ (2.3)
$_{2}^{-,*}+\textnormal{O}_{2}~{}(\textnormal{H}_{2}\textnormal{O})$ O ) $\rightarrow$ $\textnormal{O}_{2}^{-}+\textnormal{O}_{2}^{*}~{}(\textnormal{H}_{2}\textnormal{O}^{*}).$ (2.4)
Both $\mathcal{A}_{\text{gas}}$ and $v_{d}$ depend solely on the pressure-reduced electric field ($E/P$), therefore not showing separate $E$ and $P$ dependencies. [^3] This is a direct consequence of the fact that $E/P$ determines uniquely the electron energy distribution. Typical pressure-reduced drift fields in TPCs are found in the range of 10 – few 100’s V/cm/bar, which are expected to be below the dissociative-attachment thresholds for either CF4, O2 or H2O. This observation has been backed with the state-of-the-art Pyboltz transport code [24], thus anticipating BBH to be the dominant attachment mechanism in the conditions discussed in this work.
@@ 2.1 Impact of impurities on TPC response
An $E/P$ of 100 V/cm/bar would correspond to an imposing 500 kV cathode bias on a 5 m drift distance at 10 bar. Based on the studies of breakdown voltage for Ar/CF4 (99/1) carried out in [25], as well as the availability of commercial feedthroughs, we will restrict our discussion to a cathode bias not greater than 250 kV ($E/P=50$ V/cm/bar). That would roughly double the value of $\mathcal{A}_{\text{gas}}$ compared to $E/P=100$ V/cm/bar, extrapolating from [9]. Based on this, we set an O2 concentration limit of 0.1 ppm in this work to ensure charge losses remain well below 10%. [^4] Technically, the drift velocity in eq. 2.1 depends on impurities, too, as discussed in next sub-section. Fig. 2-right illustrates the impact of N2 (that is similar to that of O2 and much milder than H2O). At N2 (O2) concentrations up to 200 ppm, as the ones employed in some of the measurements presented in [9] and well above those discussed in this work, the dependence of $v_{d}$ on impurities is % level. This extrapolates to % level variations in the attachment coefficient. While these variations are negligible for attachment they can instead limit tracking severely, as discussed in text.
Modifications of the drift velocity
Impurities, even in trace-amount, can substantially modify the drift velocity to the extent of limiting particle tracking, too. Fig. 2 provides the sensitivity of Ar/CH4 (90/10) to H2O (left) and N2 (right), as a function of the pressure-reduced electric field in the range 5-50 V/cm/bar. It represents the magnitude $\left|\frac{\Delta{v_{d}}}{v_{d}}\right|=\left|\frac{v_{d}(f=0)~{}-~{}v_{d}(f)}{v_{d}(f=0)}\right|$, obtained with Pyboltz at $P=10~{}$bar and $T=25~{}^{\circ}$. The figure shows how, as expected, the impact of H2O is stronger than that of N2 by at least one order of magnitude.555 Note that achieving the sensitivity reported in the figure required using a computing cluster over few days. In case of water, this allows seeing an interesting behaviour for the 98/2 admixture, for which the $v_{d}$ plateau is within the range of fields explored. In this case, the variability increases below the plateau and, within it, it shows a compensating effect at specific fields (i) at 45 V/cm/bar for 100 ppm; ii) in the case of 1000 ppm, seemingly outside the region studied; iii) in the case of 10 ppm, statistical fluctuations limit the estimate, but a clear decrease is seen in the region 40-45 V/cm/bar). It is no surprise that the region of the plateau shows less sensitivity to impurities, as the impact of the cooling-effect of the H2O molecule in the energy distribution of the electrons mimics effectively a change in the field, that would leave $v_{d}$ unchanged. However, even in the plateau, this equivalence between cooling and field only appears to apply in certain field regions. This non-trivial effect, that could find practical application if experimentally confirmed, seems to arise from the fact that the $v_{d}$ plateau occurs when the energy distribution of the electrons is within the steep-varying Ramsauer minimum of the elastic cross-section of the main gas. This makes the product of both (energy distribution and elastic cross-section), that is the key magnitude that enters in the evaluation of $v_{d}$, very sensitive to small modifications of the shape of the energy distribution.
@@
$\Delta{z}=L\left|\frac{\Delta v_{d}}{v_{d}}\right|\simeq\frac{L}{v_{d}(f=0)}\cdot{\left|\frac{dv_{d}}{df}\right|}\cdot f,$ (2.5)
with $L=5$ m being the drift distance. The linear approximation on the r.h.s. of eq. 2.5 allows extrapolation to low concentrations ($f$), for which the numerical precision of simulations is lost. Fig. 3 presents the maximal track distortion as a function of the concentration of impurities (from eq. 2.5) for two typical electric fields (200 V/cm, 400 V/cm) and three gas mixtures (Ar/CH4 at 90/10 and 98/2, and Ar/CF4 at 99/1). For reference, the 1 mm space-resolution landmark is indicated with a horizontal line. As seen, depending on the system requirements, the tolerable impurity levels range from 10 ppm of N2 (1 mm distortion) down to values as low as 0.05 ppm of H2O (0.1 mm distortion). The case of O2 is not shown since the contamination levels that are acceptable for tracking are well above those for H2O and, in most practical conditions, its presence will impact the electron lifetime much before it starts to be a concern for tracking. Besides the aforementioned limits for O2, the tolerable concentrations of N2 and H2O that will be assumed throughout this work will be 10 ppm and 1 ppm, respectively (limited by distortions in tracking at mm-level). They are compiled in table 1.
@@
$P_{Q}=1-P_{\text{scin}}\simeq 1-\frac{1}{k\cdot P\cdot f\cdot\tau_{3}}.$ (2.6)
Here $P_{\text{scin}}$ is the scintillation probability, $k$ the quenching rate in units of s-1 bar-1 (room temperature assumed), and $f$ is the fraction of contaminant. Measurements of the scintillation quenching in [31] point to 0.2 ppm as the concentration at which CO2 would cause 10%-level light losses, and about 20 ppm for N2 (at 10 bar pressure).
Species Limit value Limiting factor
O2 $\sim$ 0.1 ppm Attachment
H2O $\sim$ 1 ppm Drift velocity
N2 $\sim$ 10 ppm Drift velocity
Table 1: Limiting values for each contaminant.
The experimental values obtained above, for the maximum concentrations of impurities that are compatible with argon scintillation, are at the same level (or even below) than those obtained from the analysis of the charge properties in regard to attachment and drift velocity. The situation for Ar/CF4 (99/1) is expected to be much more favourable, though. First, about 90% of the usable scintillation lies in the near-UV and visible range (200-800 nm) in that case [33, 16]. Considering the absorption coefficient for H2O at the low-end of such wavelength range, a value of $\Pi=10^{-2}$ cm-1 bar-1 is expected [30]. As a result, 10%-level light losses would require of H2O concentrations as high as 2000 ppm (being water the most opaque contaminant, a priori). Regarding light quenching, it has been measured in [34] for pure CF4 at 1 bar to be 50% under a N2 contamination of 4%. Assuming a quenching relation of the type 2.6, 4000 ppm of N2 would be needed to cause a 10% light loss. The factor 200 above the values obtained for pure argon matches well the expectation in eq. 2.6 resulting from the much shorter lifetime of CF4 states, in the range 10-20 ns [15]. Following the analogy with the argon case further, a plausible estimate of the maximum amount of (more reactive) molecules such as CO2 and H2O would be 40 ppm, greatly exceeding the upper bounds obtained from charged transport.